OpenStudy (anonymous):

What value of r makes the equation true ? 1/4(4r-1)=2r+1/8 A: r=-3/8 B: r= -1/6 C: r=1/4 D: 1/2

OpenStudy (anonymous):

just only subtitute

OpenStudy (anonymous):

idk how since its fractions

OpenStudy (anonymous):

First simplify left side denominator by distributive property a*(x+y) = a*x+ a*y so 4*(4r - 1) = -------------?

OpenStudy (anonymous):

16r - 4

OpenStudy (mathmale):

Could you possibly take a screen shot of the original problem and share it here?

OpenStudy (anonymous):

Yes so we have 1 /16r - 4 = 2r+1/8 Now are you familiar with cross multiplication ?

OpenStudy (anonymous):

Cross multiplication if a /b = c / d then a*d = b*c Okay ??

OpenStudy (anonymous):

Do you get it ?

OpenStudy (anonymous):

noo

OpenStudy (anonymous):



OpenStudy (anonymous):

have you understood cross multiplication method ?

OpenStudy (anonymous):

Okay or can you tell me what would be the LCD ?